(Y-2x^3) dx - x(1-xy) dy=0 find the exact equation.
 

Grade:12th pass

1 Answers

Antara
2 years ago
(y-2x3)dx=x(1-xy)dy
=> ydx-2x3dx=xdy-x2ydy
=>(xdy-ydx)/x2=ydy-2xdx
=>d(y/x)=ydy-2xdx    ….....This is the exact differential eqn.
 
Hope it helps :))